I have two problems .. please help..
Find the measure of each angle A=(x+10), b=(x-10), c=(3x-50)
and
RI invested some money at 9% and $100 less than at 7%. Her total annual interest was $81. How much did she invest at each rate?

Answer by stanbon(75887)

Find the measure of each angle A=(x+10), b=(x-10), c=(3x-50)
---
Equation:
x+10 + x-10 + 3x-50 = 180
5x-50 = 180
5x = 230
x = 46
---
A = x+10 = 56 degrees
---
B = x-10 = 36 degrees
---
C = 3x-50 = 3*46-50 = 138-50 = 88 degrees
===================================================
and
RI invested some money at 9% and $100 less than at 7%. Her total annual interest was $81. How much did she invest at each rate?
---
Interest Equation:
0.09x + 0.07(x-100) = 81
Multiply thru by 100 to get:
9x + 7x-700 = 8100
16x = 8800
x = $550 (amt. invested at 9%)
----
x-100 = $450 (amt. invested at 7%)
========
Cheers,
Stan H.
=========